# TwinMarket: A Scalable Behavioral and Social Simulation for Financial Markets

### v1

@@ 2 Background
Traditional Agent-Based Models face two major challenges in simulating complex social systems: (1) customizing sophisticated economic decision-making for individual agents and (2) enabling agents to actively perceive and respond to external environmental changes. Recent research has made progress by incorporating LLMs into agent-based simulations. CompeteAI [67] introduces a framework for modeling competitive dynamics; however, its agents rely on predefined prompts, limiting their ability to simulate nuanced economic behavior. EconAgent [41] designs economic agents with perception, reflection, and decision-making capabilities to model macroeconomic dynamics but overlooks the influence of multi-agent interactions on emergent macroeconomic phenomena. Similarly, ASFM [28], which models stock market responses to regulatory policies, fails to adequately account for the role of agent interactions and collective decision-making in shaping market trends.
In our TwinMarket framework that leverages LLM-based agents to simulate complex human behaviors, including rational decision-making, technical analysis, and behavioral biases such as herding and overconfidence [64, 24]. By grounding our model in established behavioral theories [42, 36, 54] and calibrating it with real-world data, we ensure both theoretical rigor and empirical relevance. In TwinMarket, agents make investment decisions within a socially embedded environment where interactions occur through price dynamics and social media, shaping collective market behavior via information propagation, behavioral imitation, and interaction mechanisms.
This work demonstrates the potential of LLMs to bridge computational simulations and behavioral science by systematically modeling micro-level cognitive processes and their macro-level societal effects. By integrating LLMs with established theoretical frameworks, TwinMarket provides an experimental platform to study how individual decision-making aggregates into emergent social and economic phenomena. Our findings illustrate how LLMs can simulate real-world behaviors, validate behavioral theories, and uncover mechanisms underlying social emergence, advancing our understanding of complex human systems.
@@ 3 Framework of TwinMarket
At the micro level, we focus on realistically modeling individual user behavior in the stock market, capturing decision-making processes and interactions based on real-world data. Moving to the macro level, we construct an interactive, forum-like environment where distributed agents can exchange information, form opinions, and influence each other’s decisions, thereby simulating the emergent properties of real-world social systems.
@@ 3.1 Data Sources and Simulation Setup
Using real transaction data and user profiles from Xueqiu, we computed behavioral biases following Sui’s pioneering work [56] and extracted key characteristics like demographics to construct a joint distribution. This distribution was then used to sample and generate distinct agents, each initialized with trading records to provide a historical trading baseline. Figure 2 illustrates an example of a generated agent, with further details on the generation process provided in Appendix C.
@@ 3.2 Micro-Level Simulation: Individual Behaviors
At the micro-level, our agents are equipped with three key functionalities: signal perception, planning, and decision-making. These capabilities allow them to replicate the nuanced behaviors of real-world investors. While financial decision-making may appear complex, it can be broken down into a sequence of smaller, observable actions. For example, individuals receiving market information such as stock price movements, breaking news, or social media discussions process this information to align it with their strategic objectives (desire), form an understanding of the market (belief), and ultimately execute specific trading or interact with other users through social media (intention). By replicating these granular decision-making processes, the agents establish a solid foundation for simulating the broader social structures and dynamics of financial markets.
@@ 3.3 Macro-Level Simulation: Social Interactions
In the financial market, information flow is a fundamental driver of collective behavior and market dynamics. Graphs provide a natural and powerful framework for modeling how information propagates through a network of interconnected agents, making them indispensable for macro-level simulations.
In the design of our macro-level system, we are guided by two fundamental principles. First, the graph structure must be dynamic, capable of evolving in response to the behavioral patterns of individual agents. Second, these agents should possess the ability to aggregate and process environmental information through their interconnected relationships. Our objective is to construct a framework that establishes a coherent linkage between micro-level individual interactions and macro-level systemic dynamics, thereby facilitating an adaptive and scalable architecture.
@@
$w(u,i)=\sum_{t\in T(u,i)}e^{-\lambda\cdot\Delta t},$
@@ 3.4 Workflow
To enable personalized decision-making and ensure distinct behavioral patterns across personas, we regulate each user’s perception field using recommendation algorithms. Stock recommendations are generated based on users’ portfolios and real trading data, while post scores influence stock visibility within their perception field. Once a user interacts, information propagates dynamically, forming post-propagation chains that circulate throughout the market. Before trading, users can search for relevant news and announcements (Table 1) and access technical indicators (e.g., trading volumes) and fundamental indicators (e.g., P/E ratios). To maintain valuation consistency, fundamental metrics are scaled using a ratio that aligns simulated prices with real-world values [^1] Since fundamental metrics cannot be solely derived from agent-driven trading, they are adjusted based on the relative difference between simulated and real-world prices..
@@ 4 Results and Analysis
We conduct experiments on a network of $100$ users, starting with a micro-level analysis to examine individual decision-making differences across various scenarios. We then shift to a macro-level perspective, evaluating how collective behaviors influence overall market dynamics. Our findings highlight two key aspects:
  • We observed phenomena that were previously difficult to capture using traditional ABMs, including the complete chain of events where self-fulfilling prophecies and rumors trigger collective behavior, leading to market fluctuations.
  • We explore the dynamic nature of market responses to information propagation, revealing how shifts in sentiment and information diffusion can drive fluctuations in trading behaviors and market stability.
@@ 4.3 Macro-analysis: Stylized Facts
LLM-based users inherently integrate the aforementioned behavioral traits, as LLMs have been demonstrated to analyze financial markets using both fundamental and technical metrics [16, 65]. These agents naturally exhibit herding tendencies [64]. To further capture the heterogeneity of market participants, we construct personalized user profiles for each users.
We compared the key stylized facts of TwinMarket with historical market data, ensuring that our simulations accurately reflect real-world market behaviors. Key comparisons of these stylized facts are shown in Figure 5.
  • Non-Normality: The log returns exhibit a sharp peak and fat tails.
  • Leverage Effect: Negative returns show weak autocorrelation, indicating persistence in downturns.
  • Volume-Return Relationship: Trading volume is positively correlated with returns, which typically reflects collective behavior among investors.
  • Volatility Clustering: GARCH analysis [^3] GARCH model is widely used to model time-varying volatility in financial markets. In the GARCH framework, $\alpha$ represents the sensitivity of current volatility to past returns or shocks, while $\beta$ captures the persistence of past volatility. shows that $\alpha+\beta$ is close to 1, indicating high persistence of volatility shocks.
@@ 5 Emergence of Group Behaviors
In financial markets, individual decisions are not made in isolation but are embedded within a complex social network. Price feedback, social interactions, and information flow create a dynamic feedback loop that amplifies or dampens individual actions. Through mechanisms like social influence and herd behavior, individual choices are continually reshaped by the collective actions and expectations of others, leading to the emergence of market trends and fluctuations. By precisely modeling agent micro-behaviors and their social environment, we capture group behavior patterns and emergent phenomena that align with both economic theory and real-world observations, all without relying on pre-determined rules or shortcuts.
@@ 5.1 Information Propagation
Figure 6(a) shows that beliefs in normal and rumor scenarios increasingly diverge [^4] In rumor-exposed scenarios, agents maintained 27.5% lower valuations than the control group., forming distinct echo chambers [51, 47]. This occurs as negative rumors replace normal news, shifting users’ market expectations [58]. Rumor-exposed users become more pessimistic, reassessing asset values downward, reinforcing their isolation, and further distancing themselves from unaffected groups. This process exemplifies how biased information amplifies belief divergence and strengthens echo chambers in financial markets.
@@
Trading behavior in TwinMarket is primarily driven by users’ intentions and beliefs about the market, which in turn influence their buying and selling decisions. When belief evaluations fluctuate, particularly in response to rumors, these changes significantly impact trading activity [4, 66]. As shown in Figure 6(b), under the influence of negative rumors, the proportion of sell orders increased significantly. The Sell/Buy ratio to surge to 2.02$\times$ the baseline5550.997 vs 0.495, indicating amplified panic-driven selling behavior.
@@ 5.2 Scaling Up Design
To demonstrate the robustness of our framework, we expanded our simulation to include 1,000 agents, a scale typically challenging for traditional agent-based models. The results, as depicted in Figure 9, reveal a very similar overall trend and movement throughout the five-month simulation period. This finding aligns with research showing that complex systems can exhibit similar macroscopic behavior when scaled up [7], even if the individual interactions at the micro-level become more numerous. In our case, it appears that the overall market dynamics can be well reproduced through emergent behaviors of LLM-powered agents. The technical details of our experiments, along with a more comprehensive quantitative evaluation, including metrics such as RMSE, MAE, and lag analysis, are detailed in Appendix E.2.
@@ C.3 Generating persona
For each user, we included their gender, location, and social media follower count, as these dimensions have been shown to significantly influence investor behavior [8, 25]. Additionally, we incorporated four behavioral bias indicators that we previously calculated. We sampled these user characteristics based on the joint distribution of the data in the Xueqiu dataset, thus generating the basic profiles for each user.
In the original dataset, the behavioral biases of real users were divided into three groups—“high,” “medium,” and “low”—to represent the severity of each bias. For biases with different severity levels, we created contextually realistic prompts to guide the user’s behavior accordingly. In Appendix E.1, we provide examples demonstrating how users exhibit human-like behavioral biases in their investment decisions.
For each generated user, we use the GPT-4o model to score their level of rationality. The top 40% of users, based on their rationality score, adopt a fundamental analysis investment strategy, while the remaining users follow a technical analysis strategy. Additionally, the top 10% of the most rational users are allocated ten times the initial funds compared to regular users. The distribution of strategy adoption and initial fund allocation is consistent with the findings from surveys of the A-share market.
@@ C.5 Belief Initialization
To ensure the realism of the simulation, we select the historical percentile of the BRAR sentiment index on the market’s starting date as the mean belief for the market. Based on this, we perform Gaussian sampling for each belief dimension of every Agent. Subsequently, using the belief score for each dimension, we employ GPT-4o to generate contextually realistic belief content for each Agent.
@@ E.1 Mirco-level Validation
Figure [18, 19, 20, 21, 22] show that the users we have constructed, guided by prompts, autonomously exhibit behavior biases similar to those of real-world investors during the investment decision-making process.
@@ Appendix F Prompt
Figures 23 to 34 show the overall process of our TwinMarket.

### v2

@@ 2 Background
Traditional Agent-Based Models face two major challenges in simulating complex social systems: (1) customizing sophisticated economic decision-making for individual agents and (2) enabling agents to actively perceive and respond to external environmental changes. Recent research has made progress by incorporating LLMs into agent-based simulations. CompeteAI [18] introduces a framework for modeling competitive dynamics; however, its agents rely on predefined prompts, limiting their ability to simulate nuanced economic behavior. EconAgent [19] designs economic agents with perception, reflection, and decision-making capabilities to model macroeconomic dynamics but overlooks the influence of multi-agent interactions on emergent macroeconomic phenomena. Similarly, ASFM [20], which models stock market responses to regulatory policies, fails to adequately account for the role of agent interactions and collective decision-making in shaping market trends. Table1 provides a comparison of these systems with TwinMarket, highlighting the differences in agent customization, social interactions, and simulation scale.
System Agent Customization Social Interactions Simulation Scale
CompeteAI Simple predefined prompts Competition between 2 restaurants Less than 100 agents
EconAgent Uses demographic info No interaction considered 100 agents
ASFM Two fixed strategies No interaction considered Not disclosed
TwinMarket Highly customizable agents Rich multi-agent interactions 1000 agents, scalable
Table 1: Comparison of TwinMarket with Existing Literature
In our TwinMarket framework that leverages LLM-based agents to simulate complex human behaviors, we include rational decision-making, technical analysis, and behavioral biases such as herding and overconfidence [21, 22]. By grounding our model in established behavioral theories [23, 24, 25] and calibrating it with real-world data, we ensure both theoretical rigor and empirical relevance. In TwinMarket, agents make investment decisions within a socially embedded environment where interactions occur through price dynamics and social media, shaping collective market behavior via information propagation, behavioral imitation, and interaction mechanisms.
This work demonstrates the potential of LLMs to bridge computational simulations and behavioral science by systematically modeling micro-level cognitive processes and their macro-level societal effects. By integrating LLMs with established theoretical frameworks, TwinMarket provides an experimental platform to study how individual decision-making aggregates into emergent social and economic phenomena. Our findings illustrate how LLMs can simulate real-world behaviors, validate behavioral theories, and uncover mechanisms underlying social emergence, advancing the academic understanding of complex human systems.
@@ 3 Framework of TwinMarket
At the micro level, we focus on realistically modeling individual user behavior in the stock market, capturing decision-making processes and interactions grounded in decision theories and validated by real-world trading patterns. Moving to the macro level, we construct an interactive, forum-like environment where distributed agents can exchange information, form opinions, and influence each other’s decisions, thereby simulating the emergent properties of real-world social systems.
@@ 3.1 Data Sources and Simulation Setup
Using real transaction data and user profiles from Xueqiu, we computed behavioral biases following Sui’s pioneering work [26] and extracted key characteristics like demographics to construct a joint distribution. This distribution was then used to sample and generate distinct agents, each initialized with trading records to provide a historical trading baseline. Additionally, each agent was initialized with a unique belief about the market environment and future expectations.Figure 2 illustrates an example of a generated agent, with further details on the generation process provided in Appendix C.
@@ 3.2 Micro-Level Simulation: Individual Behaviors
At the micro-level, our agents are equipped with three key functionalities: signal perception, planning, and decision-making. These capabilities allow them to replicate the nuanced behaviors of real-world investors. While financial decision-making may appear complex, it can be broken down into a sequence of smaller, observable actions. For example, individuals receiving market information such as stock price movements, breaking news, or social media discussions process this information to align it with their financial goals or aspirations (desire), form an understanding of the market (belief), and ultimately execute specific trading or interact with other users through social media (intention). By replicating these granular decision-making processes, the agents establish a solid foundation for simulating the broader social structures and dynamics of financial markets.
@@ 3.3 Macro-Level Simulation: Social Interactions
In the financial market, information flow is a fundamental driver of collective behavior and market dynamics. Graphs provide a natural and powerful framework for modeling how information propagates through a network of interconnected agents, making them indispensable for macro-level simulations [28].
The market we are designing is inherently dynamic, with agents whose behaviors evolve over time. This dynamic nature requires the system to be adaptive and scalable, capable of responding to shifts in individual agent behavior and interactions. To meet these requirements, we adhere to two fundamental principles. First, the graph structure must be dynamic, able to evolve in response to the changing patterns of individual agents. Second, the agents must be capable of aggregating and processing environmental information through their interconnected relationships. These principles guide the construction of a framework that links micro-level interactions with macro-level systemic dynamics, enabling the system to adapt and scale effectively.
@@
$w(u,i)=\sum_{t\in T(u,i)}e^{-\lambda\cdot\Delta t}$
@@ 3.4 Workflow
To enable personalized decision-making and ensure distinct behavioral patterns across personas, we regulate each user’s perception field using recommendation algorithms. The stock recommendation system is driven by users’ portfolios and real trading data. Meanwhile, the post recommendation system dynamically adjusts based on post scores, which fluctuate in real-time and influence their ranking and visibility. Additionally, each user interacts with these recommendations selectively, guided by their individual perception field. Once a user interacts, information propagates dynamically, forming post-propagation chains that circulate throughout the market. Before trading, users can search for relevant news and announcements (Table 2) and access technical indicators (e.g., trading volumes) and fundamental indicators (e.g., P/E ratios). To maintain valuation consistency, fundamental metrics are scaled using a ratio that aligns simulated prices with real-world values [^1] Since fundamental metrics cannot be solely derived from agent-driven trading, they are adjusted based on the relative difference between simulated and real-world prices..
@@ 4 Results and Analysis
We conduct experiments on a network of 100 users, starting with a micro-level analysis to examine individual decision-making differences across various scenarios. We then shift to a macro-level perspective, evaluating how collective behaviors influence overall market dynamics. Our findings highlight a key aspect: the dynamic interplay between information propagation and collective behavior, where self-fulfilling prophecies, rumors, and shifts in sentiment trigger market fluctuations. Specifically, we observe how these phenomena drive trading behaviors and impact market stability, phenomena that were previously difficult to capture using traditional ABMs.
@@ 4.3 Macro-analysis: Stylized Facts
LLM-based users inherently integrate the aforementioned behavioral traits, as LLMs have been demonstrated to analyze financial markets using both fundamental and technical metrics [38, 39]. These agents naturally exhibit herding tendencies [21]. To further capture the heterogeneity of market participants, we construct personalized user profiles for each users. These profiles enable agents to autonomously exhibit both rational investment strategies and irrational behavioral biases during decision-making, mirroring real-world investors. Further details on these dynamics are provided in Appendix E.1.
We compared the key stylized facts [^3] Stylized facts are empirical statistical patterns observed in real-world financial markets. They are commonly used to evaluate the effectiveness of simulated markets in replicating real-world dynamics[35]. of TwinMarket with historical market data, ensuring that our simulations accurately reflect real-world market behaviors. Key comparisons of these stylized facts are shown in Figure 5.
  • Non-Normality [40]: The log returns exhibit a sharp peak and fat tails.
  • Leverage Effect [41]: Negative returns show weak autocorrelation, indicating persistence in downturns.
  • Volume-Return Relationship [42]: Trading volume is positively correlated with returns, which typically reflects collective behavior among investors.
  • Volatility Clustering [43]: GARCH analysis [^4] GARCH model is widely used to model time-varying volatility in financial markets [44]. In the GARCH framework, $\alpha$ represents the sensitivity of current volatility to past returns or shocks, while $\beta$ captures the persistence of past volatility. shows that $\alpha+\beta$ is close to 1, indicating high persistence of volatility shocks.
@@ 5 Emergence of Group Behaviors
In financial markets, individual decisions are not made in isolation but are embedded within a complex social network. Price feedback, social interactions, and information flow create a dynamic feedback loop that amplifies or dampens individual actions. Through mechanisms like social influence and herd behavior, individual choices are continually reshaped by the collective actions and expectations of others, leading to the emergence of market trends and fluctuations. Through detailed modeling of individual agent behavior and their social interactions, our simulation demonstrates how macroscopic market phenomena emerge organically, validating core tenets of economic theory and mirroring empirical market observations, without imposing artificial constraints or pre-defined rules.
@@ 5.1 Information Propagation
Figure 6(a) shows that beliefs in normal and rumor scenarios increasingly diverge [^5] In rumor-exposed scenarios, agents maintained 27.5% lower valuations than the control group., forming distinct echo chambers [50, 51]. This occurs as negative rumors replace normal news, shifting users’ market expectations [52]. Rumor-exposed users become more pessimistic, reassessing asset values downward, reinforcing their isolation, and further distancing themselves from unaffected groups[53]. This process exemplifies how biased information amplifies belief divergence and strengthens echo chambers in financial markets.
@@
Trading behavior in TwinMarket is primarily driven by users’ intentions and beliefs about the market, which in turn influence their buying and selling decisions. When belief evaluations fluctuate, particularly in response to rumors, these changes significantly impact trading activity [54, 55]. As shown in Figure 6(b), under the influence of negative rumors, the proportion of sell orders increased significantly. The Sell/Buy ratio to surge to 2.02$\times$ the baseline6660.997 vs 0.495, indicating amplified panic-driven selling behavior.
@@ 5.2 Scaling Up Design
To demonstrate the scalability of our framework, we expanded our simulation to include 1,000 agents, a scale typically challenging for traditional agent-based models. The results, as depicted in Figure 9, reveal a very similar overall trend and movement throughout the five-month simulation period. This finding aligns with research showing that complex systems can exhibit similar macroscopic behavior when scaled up [57], even if the individual interactions at the micro-level become more numerous. In our case, it appears that the overall market dynamics can be well reproduced through emergent behaviors of LLM-powered agents. A more comprehensive quantitative evaluation, including metrics such as RMSE, MAE, and lag analysis, are detailed in Appendix E.2.
@@ C.3 Generating persona
For each user, we included their gender, location, and social media follower count, as these dimensions have been shown to significantly influence investor behavior [70, 74]. Additionally, we incorporated four behavioral bias indicators that we previously calculated(Table 6). We sampled these user characteristics based on the joint distribution of the data in the Xueqiu dataset, thus generating the basic profiles for each user.
In the original dataset, the behavioral biases of real users were divided into three groups—“high,” “medium,” and “low”, to represent the severity of each bias. For biases with different severity levels, we created contextually realistic prompts to guide the user’s behavior accordingly. In Appendix E.1, we provide examples demonstrating how users exhibit human-like behavioral biases in their investment decisions.
For each generated user, we use the GPT-4o model to score their level of rationality. The top 40% of users, based on their rationality score, adopt a fundamental analysis investment strategy, while the remaining users follow a technical analysis strategy. Additionally, the top 10% of the most rational users are allocated ten times the initial funds compared to regular users. The distribution of strategy adoption and initial fund allocation is consistent with the findings from surveys of the A-share market [75].
@@ C.5 Belief Initialization
To initialize the beliefs of users in a realistic manner, we first determine the mean belief for the market using the historical percentile of the BRAR sentiment index [^7] The BRAR sentiment index is a technical indicator derived from the balance of rising and falling stocks in the market, reflecting market sentiment. Data source: https://tushare.pro/. on the market’s starting date. We then sample belief values for each user across multiple dimensions using a Gaussian distribution centered around this mean. Finally, based on the sampled belief scores, we employ GPT-4 to generate contextually realistic and personalized belief content for each user, ensuring that their beliefs are both diverse and grounded in realistic market conditions.
@@ E.1 Mirco-level Validation
Figure [18, 19, 20, 21, 22, 23, 24] show that the users we have constructed, guided by prompts, autonomously exhibit both rational investment strategies and irrational behavioral biases during decision-making, mirroring real-world investors.
@@ E.2 Scaling up Validation
Today’s Trading Summary: Regarding TTEI (Technology and Telecom Index),tech growth stocks are currently at their lowest valuation in the past three years, with a PE ratio below 30. From a fundamental perspective, this sector already holds certain investment value. Although the leading tech stocks I own, such as SMIC and Haiguang Information, have been slightly affected by market sentiment in the short term, the long-term outlook remains positive, especially given the national strategy of promoting technological self-reliance. These companies’ growth potential is still promising, so today I decided to increase my position to 35%. As for CPEI (Chemical and Pharmaceutical Index), the pharmaceutical sector has undergone some adjustments, and valuations have fallen to a relatively reasonable range. My position in this sector is currently 30%, and I plan to continue holding and monitoring it. My total portfolio is now at 65%, which is relatively conservative, with the remaining 35% in cash, ready to gradually increase positions when better buying opportunities arise in the market.
Figure 18: Example of Fundamental Analysis.
Today’s Trading Summary: Seeing the recent positive news for tech stocks, I’m planning to focus more on the TTEI (Technology and Telecom Index). I already have this index in my portfolio,and the volume-price action in the last few trading days looks solid. Notably, stocks like SMIC and Haiguang Information have broken through their 5-day moving averages, and I think there’s a good chance this momentum will continue in the short term. I’m considering adding to my position in the tech sector, aiming to raise my allocation to around 30%. That said, I need to remind myself not to get too aggressive, as other sectors in my portfolio have been performing mediocrely.The Real Estate Index (REEI) has been weak lately and has hit my stop-loss level. I plan to exit and free up some capital to invest in the tech stocks.
Figure 19: Example of Technical Analysis.
@@ Appendix F Prompt
Figures 25 to 36 show the overall process of our TwinMarket.